OpenStudy (anonymous):

Assume the y varies inversely with x and y = 5 when x = 1.3. Write the indirect variation equation for the relationship. xy = 3.8 xy = 0.26 xy = 6.5

OpenStudy (anonymous):

y varies inversely with x... y = k/x use y = 5 when x = 1.3 to find k... 5 = k/1.3 5*1.3 = k So... we have y = 6.5 / x or xy = 6.5